OpenStudy (anonymous):

1. -{6x²-3xy+[7x²-8xy-(2-2y)-10y²]}+11y² the answer is x²+y²-5xy-2y-2 I need solution.

OpenStudy (anonymous):

open () \[-\left\{ 6x^{2}-3xy+\left[ 7x^{2}-8xy-2-2y-10y^{2} \right] \right\}+11y^{2}\] open [] \[-\left\{ 6x^2-3xy+7x^2-8xy-2-2y-10y^2 \right\}+11y^2\] open{} \[-6x^2+3xy+7x^2-8xy-2-2y-10y^2+11y^2\] group like terms \[-6x^2+7x^2-10y^2+11y^2+3xy-8xy-2y-2\] solve them \[x^2+y^2-5xy-2y-2\]

OpenStudy (anonymous):

@rasecciren

OpenStudy (anonymous):

i need a medal becuz i solved it completely for u :)